OpenStudy (anonymous):

Tickets to a movie cost $7.25 for adults and $5.50 for students. a group of friends purchased 8 tickets for 52.75? A. write a system of equations to represent the situation. b) How many adult tickets and student tickets were purchased.?

OpenStudy (anonymous):

a= adults s= students 7.25a +5.5s = 52.75

OpenStudy (anonymous):

1st equation: \[7.25*x + 5.50*y = 52.75\] 2nd equation: \[x+y=8\]

OpenStudy (anonymous):

where x - adults, y - students